OpenStudy (anonymous):

what's value of cos(360+12)

OpenStudy (anonymous):

it is equal to co 12

OpenStudy (dumbcow):

cos(360) = 1, sin(360) = 0 use angle sum formula cos(360+12) = cos(12)

OpenStudy (anonymous):

cos(360 + x) =cos(4*90 +12) =cos12 answer if the value of coefficient of 90 is odd then the cosine function changes to sine function.../